# IRS on orbits and the case of $SL(2,\mathbb{R})\ltimes\mathbb{R}^{2}$

### v1

@@ 2. Preliminary
Let $\lambda$ be the $\Gamma$-invariant probability measure on ${\rm Gr}_{k}(V)$. Let $G_{\lambda}=\{\tau\in GL(V)\mid\tau(\lambda)=\lambda\}$ and $I_{\lambda}=\{\tau\in GL(V)\mid\tau~{}~{}{\rm is~{}~{}trivial~{}~{}on~{}}S\}$ where $S$ is the support of $\lambda$. Then $I_{\lambda}$ is a normal algebraic subgroup of $G_{\lambda}$. Since ${\rm Gr}_{k}(V)\simeq GL(V)/H$ where $H$ is the stablizer of a $k$-dimensional subspace, by Corollary 2.6 of [5] we get that $G_{\lambda}$ is a real algebraic group and $G_{\lambda}/I_{\lambda}$ is compact. Thus, any unipotent transformation in $G_{\lambda}$ is in $I_{\lambda}$. This proves the result. ∎
@@
Let $G$ be a connected Lie group with a closed connected Lie subgroup $S$. Let $\Gamma$ be a $\sigma$-compact subgroup of Aut$(G)$ and $U$ be a subgroup of $\Gamma$ consisting of unipotent automorphisms. Suppose there is a ${\Gamma}$-invariant probability measure $\mu$ on ${\rm Sub}_{G}$ such that $\mu(\{\tau(S)\mid\tau\in\Gamma\})=1$. Then $S$ is $N$-invariant where $N$ is the smallest normal subgroup of $\Gamma$ containing $U$.
@@
Let $\mathcal{G}$ and $\mathcal{S}$ be the Lie algebras of $G$ and $S$ respectively. Let $\mathcal{O}=\{\tau(S)\mid\tau\in{\rm Aut}~{}(G)\}$. Let $X={\rm Gr}_{k}(\mathcal{G})$ be the Grassmannian manifold of $k$-dimensional subspaces of $\mathcal{G}$ where $k$ is the dimension of $\mathcal{S}$. Let $\phi\colon\mathcal{O}\to X$ be defined by $\phi(\tau(S))={\rm d}\tau(\mathcal{S})$. Then it follows from Proposition 2.2 and Lemma 2.1 that $\phi$ is continuous.
Let $\lambda=\phi(\mu)$. Consider the canonical action of $\Gamma$ on $X$ given by: $(\tau,V)\mapsto{\rm d}\tau(V)$ for $\tau\in\Gamma$ and $V\in X$. Then $\lambda$ is a $\Gamma$-invariant probability measure on $X$. By Lemma 2.4, $\lambda$ is supported on $U$-invariant subspaces. This proves that $\mathcal{S}$ is $N$-invariant. Since $S$ is a connected Lie subgroup, $S$ is $N$-nvariant. ∎
@@ 3. IRS on orbits
In this section, we show the existence and non-existence of invariant random subgroups supported on various orbits of the conjugation action of a connected Lie group $G$ on SubG. In other words, we discuss the proofs of Theorems 1.1, 1.3, 1.4 and 1.6. We also discuss certain structural results related to these theorems.
@@
It is sufficient to prove that (2) implies (3). Let $L$ be a semisimple Levi subgroup of $G$ and $N$ be the nilpotent radical of $G$ as in the statement of the theorem. Then $L$ is a connected Lie subgroup of $G$. Suppose that there is an IRS supported on the set $L_{G}$ of all semisimple Levi subgroups. Since any two semisimple Levi subgroups conjugate in $G$ by an element of $N$, there is a $N$-invariant probability measure $\mu$ such that $\mu(\{xLx^{-1}\mid x\in N\})=1$. Let $\mathcal{G}$ be the Lie algebra of $G$. Since $N$ is a nilpotent normal subgroup of $G$, Ad$(x)$ is a unipotent transformation on $\mathcal{G}$ for all $x\in N$. Therefore, by Proposition 2.5, $\mu$ is supported on $N$-invariant subspaces of $\mathcal{G}$. Thus, $L$ is normalized by $N$. Since any two semisimple Levi subgroups are conjugate by an element of $N$, we get that $L$ is normal in $G$.
Suppose the set of all semisimple Levi subgroups is closed in SubG. Then $L_{G}$ is compact; and, since $N$ is amenable, $L_{G}$ admits an $N$-invariant probability measure. Therefore, $(2)$ implies $(3)$ proves that $L$ is normal in $G$.
@@
Now, let $N$ be the nilpotent radical of $G$. Then Ad$(x)$ is unipotent for any $x\in N$, hence $N$ normalizes $K$. Thus, $NK$ is a closed connected subgroup of $G$. Since $K$ is a maximal compact subgroup of $G$, $K$ is a maximal compact subgroup of $NK$. Since $N$ normalizes $K$, $K$ is normal in $NK$ and hence contains any compact subgroup of $NK$. Recall that any two maximal compact subgroups of a connected Lie groups are conjugate. Let $T=N\cap K$. Then $T$ is the largest compact normal subgroup of $N$ and hence $N/T$ is a simply connected nilpotent Lie group. Since $N$ is normal in $G$ and $T$ is the largest compact normal subgroup of $N$, $T$ is normal in $G$. Since $T\subset K$, replacing $G$ by $G/T$, we may assume that $N$ is simply connected.
@@
Let $G$ be a connected Lie group with solvable radical $R$ and $M$ be a maximal compact subgroup of $G$. Suppose $M$ is normal in $G$. Then we have the following:
@@
    solvable radical $R$ admits a compact central subgroup $T$ so that $R/T$ is simply connected;
@@
Since $B$ is a Borel subgroup of $G$, $G/B$ is compact, and hence there is a compact set $C$ in $G$ such that $G=CB$. Since any two Borel subgroups are conjugate in $G$, $B_{G}=\{xBx^{-1}\mid x\in C\}$. Thus, $B_{G}$ is a compact subset of SubG.
@@
Suppose that there is an IRS $\mu$ such that $\mu(D_{G})=1$. Then, by Proposition 2.5, $D$ is normalized by any unipotent subgroup of $G$. If $S$ is a noncompact simple algebraic subgroup of $G$, then maximal split diagonalizable subgroup $D_{S}$ of $S$ is contained in a conjugate of $D$. Replacing $S$ by a conjugate of $S$, we may assume that $D_{S}\subset D$. Since $S$ is noncompact, $S$ is generated by unipotent elements, hence $S$ normalizes $D$. Therefore, $D\cap S$ is a normal subgroup of positive dimension which is a contradiction to the simplicity of $S$. Thus, the semisimple Levi subgroup $L^{\prime}=[L,L]$ of $G$ is compact. This implies that $D\subset Z(L)$ because $L=[L,L]Z(L)$.
@@
Since $D$ is normalized by the unipotent subgroups, $D$ is normalized by $U$. Since $D\cap U$ is trivial, $D$ is centralized by $U$ also. Since $G=LU$, $D$ is central in $G$.
The maximal diagonalizable subgroup being normal has certain implications on the structure of $G$ which is noted down below.
Proposition 3.3.
Let $G$ be a connected real algebraic group and $D$ be a maximal diagonalizable subgroup of $G$. In addition, suppose $D$ is normal in $G$. Then any Levi subgroup of $G$ is a compact subgroup, and the solvable radical is a compact extension of a nilpotent group.
Proof.
Let $H$ be a noncompact simple factor of a semisimple Levi subgroup of $G$ and $D_{H}$ be the maximal diagonalizable subgroup of $H$. Then $D_{H}$ is contained in a conjugate of $D$. Since $D$ is normal, $D_{H}\subset D$. This implies that $D\cap H$ is an abelian normal subgroup of positive dimension in $H$, which is a contradiction to the simplicity of $H$. Therefore, any semisimple Levi subgroup of $G$ is compact.
@@
Now, let $R$ be the solvable radical of $G$. Then $R\cap D$ is a normal subgroup $R$, hence $D_{R}U$ is a nilpotent subgroup of $R$ where $D_{R}=D\cap R$ and $U$ is the unipotent radical of $G$. Since $R/D_{R}U$ is a anisotropic torus (that is, without any split torus subgroup) we get that $R/D_{R}U$ is compact. ∎
4. Application to measure preserving actions
Let $G$ be a locally compact group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. It follows from [14] that the stabilizer of almost every point in X is closed in G and the stabilizer map $x\mapsto G_{x}$ is measurable where $G_{x}=\{g\in G\mid gx=x\}$ for $x\in X$. Thus, the $G$-invariant probability measure $m$ on $X$ may be realized as an IRS on G which leads to the following:
Corollary 4.1.
Let $G$ be a connected Lie group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. Then almost none of the stabilizer subgroups is a semisimple Levi subgroup (resp. maximal compact subgroup) of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a semisimple Levi subgroup (resp. maximal compact subgroup) of $G$ and $G_{x}$ is normal in $G$. In addition, if the action of $G$ is ergodic, then almost none of the stabilizer is a semisimple Levi subgroup of $G$ or the action of $G$ factors through the solvable radical of $G$.
Proof.
Let $L_{G}$ be the set of all semisimple Levi subgroups of $G$ and $X_{L}=\{x\in X\mid G_{x}\in L_{G}\}$. Since $L_{G}$ is invariant under the conjugation action and $G_{gx}=gG_{x}g^{-1}$, it follows that $X_{L}$ is invariant under the action of $G$. Suppose $m(X_{L})>0$. Then, replacing $X$ by $X_{L}$ and $m$ by the restriction of $m$ to $X_{L}$, we obtain a measure preserving action of $G$ such that almost all stabilizers are Levi subgroups of $G$. Using the identification of $x\mapsto G_{x}$, we obtain a IRS supported on the set of semisimple Levi subgroups. By Theorem 1.1, we get that $G$ has a unique semisimple Levi subgroup, say L, and $L$ is normal in $G$. The case of maximal compact subgroups may be proved in a similar way.
Suppose the action of $G$ is ergodic. Then $X_{L}$ has full measure or null. In case $X_{L}$ has full measure, any $G_{x}=G_{y}$ for all $x,y\in X_{L}$ and hence the semisimple Levi subgroup acts trivially on $X$, hence the action factors through the solvable radical of $G$.
A similar application can also be done for maximal diagonalizable groups and Borel subgroups as in the following: we omit the details as it is same as in Corollary 4.1
Corollary 4.2.
Let $G$ be a connected real algebraic group and $(X,m)$ be a probability space on which $G$ acts by measure preserving transformations. Then almost none of the stabilizer subgroup is a Borel (resp. maximal diagonalizable) subgroup of $G$ or $m$ has an atom $x$ such that $G_{x}$ is a Borel (resp. maximal diagonalizable) subgroup of $G$ and $G_{x}$ is normal in $G$. In addition if the action of $G$ is ergodic, then almost none of the stabilizer is a Borel (resp. maximal diagonalizable) subgroup of $G$ or the action of $G$ factors through a compact subgroup (resp. compact extension of the unipotent radical) of $G$.
Corollary 4.3.
Let $G$ be a connected Lie group and $K$ be a maximal compact subgroup of $G$. Suppose $G/K$ carries a $G$-invariant probability measure. Then $G$ is a compact group.
Proof.
Let $m$ be the $G$-invariant probability measure on $G/K$. Then $(G/K,m)$ is a probability space on which $G$ acts by measure preserving transformations and stabilizer at any $xK\in G/K$ is $xKx^{-1}$. Therefore by Corollary 4.1, $K$ is normal in $G$, hence $G/K$ is a locally compact group. Since $G/K$ carries a $G$-invariant probability measure, $G/K$ is compact. Thus, $G$ is compact. ∎

### v2

@@ 3. On the support of random subgroups invariant under unipotent transformations
Let $\lambda$ be the $\Gamma$-invariant probability measure on ${\rm Gr}_{k}(V)$. Let $G_{\lambda}=\{\tau\in GL(V)\mid\tau(\lambda)=\lambda\}$ and $I_{\lambda}=\{\tau\in GL(V)\mid\tau~~{\rm is~~trivial~~on~}S\}$ where $S$ is the support of $\lambda$. Then $I_{\lambda}$ is a normal algebraic subgroup of $G_{\lambda}$. Since ${\rm Gr}_{k}(V)\simeq GL(V)/H$ where $H$ is the stabilizer of a $k$-dimensional subspace, by Corollary 2.6 of [7] we get that $G_{\lambda}$ is a real algebraic group and $G_{\lambda}/I_{\lambda}$ is compact. Thus, any unipotent transformation in $G_{\lambda}$ is in $I_{\lambda}$. This proves the result. ∎
The following proposition is crucial to the proof of our results, it describes the action of the unipotent group on orbits that support invariant measures: for a connected Lie group $G$, Aut(G)(G) has a natural action on the Lie algebra of $G$ by identifying each $\tau\in{\rm Aut}(G)$ with its differential ${\rm d}\tau$.
@@
Let $G$ be a connected Lie group with Lie algebra $\mathcal{G}$}, and $S$ be a closed connected subgroup of $G$ with Lie algebra $\mathcal{S}$. Let $\Gamma$ be a $\sigma$-compact subgroup of Aut(G)(G), and $U$ be a subgroup of $\Gamma$ consisting of unipotent automorphisms. If either
(1) there is a $\Gamma$}-invariant probability measure $\lambda$ on ${\rm Sub}_{\mathcal{G}}$ with $\lambda(\{{\rm d}\tau({\mathcal{S}})\mid\tau\in\Gamma\})=1$
or
(2) there is a $\Gamma$}-invariant probability measure $\mu$ on ${\rm Sub}_{G}$ with $\mu(\{\tau(S)\mid\tau\in\Gamma\})=1$;
then $S$ is $N$-invariant where $N$ is the smallest normal subgroup of $\Gamma$ containing $U$
@@
Let $X={\rm Gr}_{k}(\mathcal{G})$ be the Grassmannian manifold of $k$-dimensional subspaces of $\mathcal{G}$ where $k$ is the dimension of $\mathcal{S}$. Consider the canonical action of $\Gamma$ on $X$ given by: $(\tau,V)\mapsto{\rm d}\tau(V)$ for $\tau\in\Gamma$ and $V\in X$.
If (1) holds, then in view of Lemma 2.1, we get that $\lambda$ is a $\Gamma$-invariant probability measure on $X$. By Lemma 3.1, $\lambda$ is supported on $U$-invariant subspaces. This proves that $\mathcal{S}$ is NN-invariant. Since $S$ is a connected Lie subgroup, $S$ is NN-nvariant.
Suppose (2) holds. Let $\mathcal{O}=\{\tau(S)\mid\tau\in{\rm Aut}~(G)\}$ and $\phi\colon\mathcal{O}\to X$ be defined by $\phi(\tau(S))={\rm d}\tau(\mathcal{S})$. Then it follows from Proposition 2.2 and Lemma 2.1 that $\phi$ is continuous. Let $\lambda=\phi(\mu)$. Then $\lambda$ is a $\Gamma$-invariant probability measure and $\lambda(\{\tau({\mathcal{S}})\mid\tau\in\Gamma\})=1$. Now the result follows from the previous case. ∎
@@ 4. IRS on orbits
In this section, we show the existence of invariant random subgroups supported on various orbits of the conjugation action of a connected Lie group $G$ on $\text{Sub}_{G}$. In other words, we discuss the proofs of Theorems 1.1, 1.3, 1.4 and 1.6. We also discuss certain structural results related to these theorems.
@@
It is sufficient to prove that (2) implies (3). Let $L$ be a semisimple Levi subgroup of $G$ and $N$ be the nilpotent radical of $G$ as in the statement of the theorem. Then $L$ is a connected Lie subgroup of $G$. Suppose that there is an IRS supported on the set $L_{G}$ of all semisimple Levi subgroups. Since any two semisimple Levi subgroups conjugate in $G$ by an element of $N$, there is a $N$-invariant probability measure $\mu$ such that $\mu(\{xLx^{-1}\mid x\in N\})=1$. Let $\mathcal{G}$ be the Lie algebra of $G$. Since $N$ is a nilpotent normal subgroup of $G$, Ad(x)(x) is a unipotent transformation on $\mathcal{G}$ for all $x\in N$. Therefore, by Proposition 3.2, $\mu$ is supported on $N$-invariant subspaces of $\mathcal{G}$. Thus, $L$ is normalized by $N$. Since any two semisimple Levi subgroups are conjugate by an element of $N$, we get that $L$ is normal in $G$.
Suppose the set of all semisimple Levi subgroups is closed in $\text{Sub}_{G}$. Then $L_{G}$ is compact; and, since $N$ is amenable, $L_{G}$ admits an $N$-invariant probability measure. Therefore, $(2)$ implies $(3)$ proves that $L$ is normal in $G$.
@@
Now, let $N$ be the nilpotent radical of $G$. Then Ad(x)(x) is unipotent for any $x\in N$, hence $N$ normalizes $K$. Thus, $NK$ is a closed connected subgroup of $G$. Since $K$ is a maximal compact subgroup of $G$, $K$ is a maximal compact subgroup of $NK$. Since $N$ normalizes $K$, $K$ is normal in $NK$ and hence contains any compact subgroup of $NK$. Recall that any two maximal compact subgroups of a connected Lie groups are conjugate. Let $T=N\cap K$. Then $T$ is the largest compact normal subgroup of $N$ and hence $N/T$ is a simply connected nilpotent Lie group. Since $N$ is normal in $G$ and $T$ is the largest compact normal subgroup of $N$, $T$ is normal in $G$. Since $T\subset K$, replacing $G$ by $G/T$, we may assume that $N$ is simply connected.
@@
Let $G$ be a connected Lie group with solvable radical $R$ and $M$ be a maximal compact subgroup of $G$. Suppose $M$ is normal in $G$. Then,
@@
    the solvable radical $R$ admits a compact central subgroup $T$ so that $R/T$ is simply connected;
@@
We now apply the above results to show that a finite covolume subgroup is cocompact if the subgroup is a Levi subgroup or a maximal compact subgroup.
Corollary 4.3.
Let $G$ be a connected Lie group and $H$ be a closed subgroup of $G$ such that $G/H$ carries a $G$-invariant probability measure. Suppose $H$ is either a Levi subgroup or a maximal compact subgroup. Then $G/H$ is a compact group.
Proof.
We prove the result when $H$ is a Levi subgroup, the other case is similar. Suppose $H$ is a Levi subgroup of $G$. Then the stabilizer, say $G_{xH}$, at any $xH\in G/H$ is $xHx^{-1}$. Therefore, using the map $xH\mapsto G_{xH}$, we obtain a IRS supported on the set of semisimple Levi subgroups. By Theorem 1.1, we get that $G$ has a unique semisimple Levi subgroup and is normal in $G$. This implies that $G/H$ is a locally compact group. Since $G/H$ carries a $G$-invariant probability measure, $G/H$ is compact. ∎
@@
Since $B$ is a Borel subgroup of $G$, $G/B$ is compact, and hence there is a compact set $C$ in $G$ such that $G=CB$. Since any two Borel subgroups are conjugate in $G$, $B_{G}=\{xBx^{-1}\mid x\in C\}$. Thus, $B_{G}$ is a compact subset of $\text{Sub}_{G}$.
@@
Suppose that there is an IRS $\mu$ such that $\mu(D_{G})=1$. Then, by Proposition 3.2, Any unipotent subgroup of $G$ normalizes $D$. If $S$ is a noncompact simple algebraic subgroup of $G$, then the maximal split diagonalizable subgroup $D_{S}$ of $S$ is contained in a conjugate of $D$. Replacing $S$ by a conjugate of $S$, we may assume that $D_{S}\subset D$. Since $S$ is noncompact, $S$ is generated by unipotent elements, hence $S$ normalizes $D$. Therefore, $D\cap S$ is a normal subgroup of positive dimension which is a contradiction to the simplicity of $S$. Thus, the semisimple Levi subgroup $L^{\prime}=[L,L]$ of $G$ is compact. This implies that $D\subset Z(L)$ because $L=[L,L]Z(L)$.
@@
Since the unipotent subgroups normalize $D$, $D$ is normalized by $U$. Since $D\cap U$ is trivial, $D$ is centralized by $U$ also. Since $G=LU$, $D$ is central in $G$. This proves that (1) implies (2) and that (2) implies (3) is trivial. If $D$ is normal, the $\delta_{D}$, the dirac measure at $D$ is the required IRS. This proves that (3) implies (1).
@@
Suppose (1) holds. Then the proof of the first implication shows that Levi subgroup is compact and $D$ is central. Let $R$ be the solvable radical of $G$. Then $D_{R}=R\cap D$ is a central subgroup of $R$, hence $D_{R}U$ is a nilpotent subgroup of $R$ where $U$ is the unipotent radical of $G$. Since $R/D_{R}U$ is an anisotropic torus (that is, without any split torus subgroup) we get that $R/D_{R}U$ is compact. ∎